OpenStudy (anonymous):

the organizers of a festival wish to sell polo shirts and sweaters with the festival logo. if x represents the number of polo shirts and y represents the number of sweaters write a 2 variable first degree inequality for each of the following statements: 1.a least 80 polo shirts 2.a maximum of 120 sweaters 3.at least 3 times as many polo shirts as sweater 4.at most as many sweaters as polo shirts

OpenStudy (anonymous):

i got most of answers just need 3 and maybe 4?. \[x \ge 80y\] \[x \ge120y\] \[3x \ge y\] \[y \le \] can someone verify please

OpenStudy (anonymous):

anyone?:P